OpenStudy (anonymous):

charge Q is distributed uniformly along a semicircle of radius a. which formula below gives correct magnitude of electric field at centre of circle?

OpenStudy (vincent-lyon.fr):

See my attachment in: http://openstudy.com/users/goutham1995#/updates/4fae5db2e4b059b524f9a07e

OpenStudy (anonymous):

\[1)E = kQ/(pi) a \] \[2. E = kQ/pi a ^{2}\] \[3. E= K * 2Q/pi*a\] 4.E=k*2Q/pie * a^2 5.E=k*2Q/a^2

OpenStudy (anonymous):

these are the options

OpenStudy (vincent-lyon.fr):

Can you find which ones to rule out because of lack of homogeneity?

OpenStudy (anonymous):

yeah good idea!

OpenStudy (anonymous):

but it will take a long time

OpenStudy (vincent-lyon.fr):

How? electric field is Coulomb's constant times charge over distance squared

OpenStudy (anonymous):

yeah but we have to do that for all the options

OpenStudy (vincent-lyon.fr):

It only takes a few seconds to see that 1) and 3) cannot be expressions for an electric field.

OpenStudy (anonymous):

yeah..because if 1 is not possible, then 3 too is not possible since it only has a constant along with it

OpenStudy (anonymous):

other three are matching

OpenStudy (vincent-lyon.fr):

But look at the denominator!

OpenStudy (vincent-lyon.fr):

There MUST be a distance squared in the denominator. If not, no chance to end up with an electric field

OpenStudy (anonymous):

yeah i understood that..hence option 2 , 4 and 5 are possible

OpenStudy (vincent-lyon.fr):

Yes. Now, 2) is not possible either because fields add up as vectors, not as scalars.

OpenStudy (anonymous):

i didnt understand that ..:(

OpenStudy (anonymous):

????????

OpenStudy (vincent-lyon.fr):

The method to derive the field is here: study each of the steps. They have numbers so it makes it easier to say later which needs elaboration.

OpenStudy (anonymous):

yeah understood the derivation. but i didnt understand how you eliminated option 2

OpenStudy (vincent-lyon.fr):

As all charges are at the same distance a of centre, then the answer given looks like a sum of magnitudes. It is not so important.

OpenStudy (anonymous):

oh it looks like k*lambda

OpenStudy (vincent-lyon.fr):

So do you know which answer is good now?

OpenStudy (anonymous):

option d or e

OpenStudy (vincent-lyon.fr):

I mean, you've just said you understand the derivation; so you must use the result to find which is the right expression.

OpenStudy (vincent-lyon.fr):

The result is in box 4. How does it apply to your problem?

OpenStudy (anonymous):

\[2*K*\lambda/a [\sin]\left(\begin{matrix}90 \\ 0\end{matrix}\right)\]

OpenStudy (vincent-lyon.fr):

Yes, so the final part is just sin90°-sin0° = 1 and you have to work out λ in terms of Q and a.

OpenStudy (anonymous):

okay thanks